# Dirac Operators on Configuration Spaces and Yang-Mills Quantum Field Theory

### v2

@@ Acknowledgements
JMG would like to express his sincere gratitude to entrepreneur Kasper Bloch Gevaldig for his continued financial support. JMG is also indepted to the following list of sponsors for their generous support: Frank Jumppanen Andersen, Bart De Boeck, Simon Chislett, Jos van Egmond, Trevor Elkington, Jos Gubbels, Claus Hansen, David Hershberger, Ilyas Khan, Simon Kitson, Hans-Jørgen Mogensen, Stephan Mühlstrasser, Bert Petersen, Ben Tesch, Jeppe Trautner, Vladimir Zakharov, and the company Providential Stuff LLC. JMG would also like to express his gratitude to the Institute of Analysis at the Gottfried Wilhelm Leibniz University in Hannover, Germany, for kind hospitality during numerous visits.

### v3

@@ Acknowledgements
JMG would like to express his gratitude to entrepreneur Kasper Bloch Gevaldig for his unwavering financial support. JMG is also indepted to Master of Science in Engineering Vladimir Zakharov and to the following list of sponsors for their generous support: Frank Jumppanen Andersen, Bart De Boeck, Simon Chislett, Jos van Egmond, Trevor Elkington, Jos Gubbels, Claus Hansen, David Hershberger, Ilyas Khan, Simon Kitson, Hans-Jørgen Mogensen, Stephan Mühlstrasser, Bert Petersen, Ben Tesch, Jeppe Trautner, and the company Providential Stuff LLC. JMG would also like to express his gratitude to the Institute of Analysis at the Gottfried Wilhelm Leibniz University in Hannover, Germany, for kind hospitality during numerous visits.